Hayhayz (hayhayz):

Where does the Pythagorean Identity sin2 Θ + cos2 Θ = 1 come from? How does it relate to right triangles?

OpenStudy (abdullahm):

 sin^2 + cos^2 = 1 a^2/c^2 + b^2/c^2 = 1 (a^2 + b^2)/c^2 = 1 c^2/c^2 = 1 1 = 1

OpenStudy (ijlal):

another way of doing is by using the formula of equation of a unit circle  for a unit circle r=1 and u can use the method of @AbdullahM by expressing x and y in terms of sine and cosine :)